# Canonicalization for Unreproducible Builds in Java

### v1

@@ 4.2. Methodology
We have a dataset of 12,283 unreproducible artifacts and for each of them we have the reference artifact from Maven Central and its rebuilt version from our own rebuild. We run diffoscope 285 on each pair of reference and rebuild artifacts to get the content diff. diffoscope is a state-of-the-art tool for computing the difference between various types of files, archives, and directories  [^10] https://diffoscope.org/. This gives us a total of 12,283 diffoscope files. The median size of diffoscope files is 102 KB, the maximum is 3.40 GB, and the minimum is 0.38 KB.
Given the scale of the dataset, we first analyze a random sample of the diffoscope files to understand the reasons of unreproducibility. There are two attributes in the diffoscope file that are of interest to us - source1 and unified_diff, both of them indicating a reproducibility problem. The source1 attribute either contains the name of the file or the command that is run to get the textual output of the file. For example, if source1 includes cyclonedx.json, we can infer that the difference is in the Software Bill of Materials (SBOM). Another example is if source1 is javap -verbose -constants -s -l -private {}, it means that the Java class file is being compared and the reason of unreproducibility is due to differences in the bytecode. The unified\_diff attribute contains the diff of the file in the unified format, for textual files. We then write regular expressions to match these attributes to categories. For example, if regex
@@
is matched, we infer that the reason of unreproducibility is due to timestamps in the file. We define regular expressions, which we match against all 12,283 diffoscope files.
Finally, we manually go through samples that remain unmatched to the regular expressions and create expressions to map them too. This refines the different root causes of unreproducibility. Then, we report on mitigation strategies from the literature that could fix the reproducibility problem.
@@ 4.3. Results
Reason for Unreproducibility Root Cause of Unreproducibility Novelty Example Main Mitigation
Build Manifests Environment - Built-By attribute Canonicalization by rebuilder
Rebuild Process - Build-Jdk attribute Fix rebuild process
Non-deterministic configuration Embedded branch names Fix build process
SBOM Java Vendor Different checksum algorithms Fix rebuild process
Custom release configuration Releasing a subset of artifacts Fix build process
Non-deterministic information Serial Number Canonicalization by rebuilder
Filesystem Environment - Permissions Canonicalization by rebuilder
Custom release configuration Generated binaries are not included Fix rebuild process
JVM Bytecode JDK Version - Ordering of constant pool entries Fix rebuild process
Embedded data Git branch embedded Fix build process
Build time generated code - Lambda functions Canonicalization by rebuilder
Versioning Properties - - Number of tags embedded in Jar Canonicalization by rebuilder
Timestamps - - Embedded in shell script in Jar files Fix build process
Table 1. The taxonomy of unreproducibility causes based on the analysis of 12,283 unreproducible artifacts. We are the first to state the most appropriate mitigation strategy.
In this section, we present a comprehensive taxonomy of unreproducibility in Java, based on our large-scale, systematic protocol presented in Figure 1. Table 1 summarizes the taxonomy, before deep-diving into each of them. The first column lists the reason for unreproducibility. The second column further breaks down the reason into root causes. The third column indicates if the root cause is novel and has not been reported in the literature before, to our knowledge. The fourth column gives an example of the root cause that appears in our dataset. Finally, the last column lists the main mitigation strategy to fix the unreproducibility, one of the three presented in subsection 2.3.
For example, consider the row about JVM bytecode, where unreproducibility is due to changes in the JVM bytecode between the reference and rebuild artifacts. One of its root cause is due to embedded data in the bytecode, this cause has not been discussed before. The main mitigation strategy for this problem is to fix the build process by preventing build tools to embed branch names in the JVM bytecode. Branch names are mutable and can change over time. Hence embedding them in the bytecode can lead to difference in subsequent builds. The other two mitigation strategy would also work, but they have cons - 1) third-party rebuilder would have to debug the build process to understand what build tool is responsible for embedding and 2) branch names have no conventional names and hence canonicalization would risk removing meaningful differences in the bytecode.
Next, We describe each reason and its root cause in details in the following subsections.
@@ 4.3.1. Unreproducible Build Manifests
We consider two manifest files in our analysis - MANIFEST.MF and pom.properties. The MANIFEST.MF file is a mandatory metadata file in a JAR file. It potentially contains information on the entry point of the JAR file, signatures of the classes, and details of the version control system  [^11] https://docs.oracle.com/javase/tutorial/deployment/jar/manifestindex.html. pom.properties is an automatically generated file by Maven that contains the group ID, artifact ID, and version of the Maven module.
The first set of differences in manifests is due to a change in environment. For example, the Built-By attribute in the MANIFEST.MF file is sometimes updated to the username of the user who built the JAR file. Mitigation: This non-deterministic information should be stripped out from the JAR file by rebuilder before comparison. See section A.1, section A.4, section A.5, and section B.3 for more attributes that follow the same problem and solution.
The next set of differences in manifests is due to differences in the script that configures the rebuild. For example, the Build-Jdk attribute depends upon the version of JDK used to build the JAR file. Mitigation: All environment specifications must be exactly specified in the the build script by rebuilder. Ideally, the build script file extracts information from the MANIFEST to be reproduced. See section A.2, section A.3, section A.6, section A.7, and section A.8 for concrete examples.
Finally, there are reasons of differences in the manifests MANIFEST.MF and pom.properties that are due to non-deterministic behavior in the build script of the builder. For example, build manifests pom.properties and MANIFEST.MF can differ in order of the attributes if they are not sorted. Mitigation: The builder should depend upon on plugins that generate sorted data  [^12] https://github.com/apache/maven-archiver/commit/763a940540eefad74f9ba73cb5eed288dc4e639d, should not embed non-deterministic data (eg branch names) in the artifact, should ignore timestamp, and stay up to date with the latest version of plugins. section A.9, section A.10, section A.11, section B.1, and section B.2 describe these reasons and solutions in detail.
@@ 4.3.2. Unreproducible Software Bill of Materials
The first set of changes are due to incorrect Java vendor used by the rebuilder. For example, old JDK versions do not have newer hash algorithms like SHA-3. However, vendors of some JDKs backport these algorithms and if builder uses a JDK with backported algorithms, the resulting SBOM will have hashes computed using more algorithms. If rebuilder does not use the same JDK, there will be differences in number of hash algorithms used. Mitigation: The rebuilder should use the same JDK vendor and version as the builder. See section C.1 for more details.
The next set of changes are due to a tailored release configuration by the builder. For example, the builder can choose to release only a subset of the artifacts produced by the build process which can lead to less components in the released SBOM. Mitigation: The builder should use a conventional build process - which is to streamline the build process only using Maven commands. This helps the rebuilder to reproduce the build process exactly without knowing all ad hoc scripting in the build. See section C.2 and section C.3 for concrete examples.
Finally, the last set of changes are due to non-deterministic information in the SBOM. CycloneDX attributes like timestamp and serialNumber are non-deterministic and hence they differ in each build. Mitigation: There are two possible mitigations. First, the builder can generate the values of these attributes so that they are deterministic. Second, the rebuilder can strip out these attributes from the SBOM before comparing the outputs. See section C.5, section C.4, and section C.6 which outlines the implementation on how to either generate deterministic values or strip out the non-deterministic values.
Reproducible SBOMs is an open problem in the field, we are the first to report about it.
@@ 4.3.3. Unreproducible Filesystem
We observe unreproducible Maven releases due to presence or absence of files, changes in the absolute file paths and their related metadata such as permissions, ownership, size and type. We classify them into 2 reasons below.
The first set of changes are due to an influence from the environment. For example, the permissions of the file can vary depending on the umask of the environment of rebuilder section D.2. The ownership and timestamp of the file can also vary depending who and when the artifact is built section D.3. File sizes can vary depending on the file system used to build the artifact section D.6. The names of the files can vary if absolute paths are used to refer to the files in artifact section D.7. These names can be of generated files or they could be embedded in JVM bytecode and configuration files which makes the artifact unreproducible section D.8. Finally, the order of files in Jar file section G.5 can vary as the order depends upon the filesystem layout  [^14] https://reproducible-builds.org/docs/archives/. Mitigation: The build script should either set a fixed value or strip this information from reference and rebuild artifacts before comparison.
Finally, there are some causes due to differences in the build process by the builder and rebuilder. For example, some binaries are not included in the rebuild version as they are not generated by the rebuilder build process. Mitigation: The rebuilder process should replicate the build process of the builder exactly. For the above example, the rebuilder should incorporate commands to generate the missing binaries. See section D.1 and section D.5 for more concrete examples.
@@ 4.3.4. Unreproducible JVM bytecode
We observe 898 unreproducible Maven releases due to changes in the JVM bytecode when rebuilding. The reasons are as follows.
@@
The first set of changes are due to the use of different versions of the JDK or JDK flags used in the build and rebuild process. This differences are due to core compilation changes such as refactoring, ordering changes in the bytecode (specifically methods, fields, static initializers, entries in the constant pool, and array type values in annotation section G.1), optimization flags, debug flags. For example, java.nio.ByteBuffer.flip() returns a different type in JDK 8 and JDK 11. For io.dropwizard.metrics:metrics-collectd:4.1.20 rebuilt with JDK 11, all flip calls return java.nio.ByteBuffer which is different from the reference version built with JDK 8 where the return type is java.nio.Buffer.  [^15] This is reported in the bug tracking system of OpenJDK (Gafter, 2002) which shows that this change is introduced in JDK 9. See section E.1, section E.2, and section E.3 for more examples. Mitigation: The rebuilder should use the same JDK version and flags as the builder. We propose that the builder should always embed the precise vendor and version of JDK used to build the artifact in the MANIFEST file under the attribute Build-Jdk. This way, the rebuilder can use the same version and vendor of JDK to rebuild the artifact by reading the information instead of guessing it. If the builder is using a custom JDK as done in Google Guava (Moore, 2023), this should be made available to the rebuilder for verification.
@@
The next set of changes are due to embedding information related to environment. We observe that the bytecode has git branch names, user name of the builder, and absolute paths. For example, org.apache.hive:hive-exec:4.0.0-alpha-2 embeds git properties and timestamps in package-info.class file using a shell script that adds a Java annotation to package-info.java. The Java compiler also embeds the exact Java version used to compile the source code in the bytecode of module-info. However, this hinders reproducibility even if the patch version of the JDK is different. Mitigation: The builder should avoid embedding environment specific information in the bytecode. This information changes with each rebuilder and hence the bytecode will be different making the artifact unreproducible. See section E.4 for example of a release where it happens.
@@
Finally, the last set of changes are due to build time code generation or modification. The Java compiler produces synthetic code during build in order to handle access to type members (Rimenti, 2018). The compiler also names lambda functions (Evans, 2020) and anonymous classes (Maverik, 2016) which are not present in the source code. In addition to these capabilities, Maven provides plugins to generate or modify code during the build process. All of this generated code can be different across different builds and hence the build is unreproducible. For example, we observe a difference in names of lambda function in org.apache.helix:zookeeper-api:1.0.3. Another example is org.apache.hive:kafka-handler:4.0.0-alpha-2 which uses maven-shade-plugin to rename package names from kafkaesqueesque to kafkaesque but only when a Maven profile is activated. Mitigation: The names of the synthetic code can be canonicalized by the rebuilder before comparison. Bytecode canonicalization (Sharma et al., 2024) is a technique to strip out non-deterministic information from the bytecode. Differences caused by Maven plugins can be mitigated by running the same plugins as in the original build process. Maven plugins can be configured to run only for certain build profiles. The rebuilder should document the build profiles used by the builder to reproduce the build process. See section E.5 for more examples.

### v2

@@ 4.2. Methodology
We have a dataset of 12,803 unreproducible artifacts and for each of them we have the reference artifact from Maven Central and its rebuilt version from our own rebuild. We run diffoscope 285 on each pair of reference and rebuild artifacts to get the content diff. diffoscope is a state-of-the-art tool for computing the difference between various types of files, archives, and directories  [^12] https://diffoscope.org/. This gives us a total of 12,803 diffoscope files.
Given the scale of the dataset, we first analyze a random sample of the diffoscope files to understand the reasons of unreproducibility. There are two attributes in the diffoscope file that are of interest to us — source1 and unified_diff, both of them indicating a reproducibility problem. The source1 attribute either contains the name of the file or the command that is run to get the textual output of the file. For example, if source1 includes cyclonedx.json, we can infer that the difference is in the Software Bill of Materials (SBOM). Another example is if source1 is javap -verbose -constants -s -l -private {}, it means that the Java class file is being compared and the reason of unreproducibility is due to differences in the bytecode. The unified_diff attribute contains the diff of the file in the unified format, for textual files. Next, we notice that the changes are present in the most deeply nested unified_diff attribute which represents the textual difference in the file extracted from the archive. This is because Diffoscope recursively extracts files from the archive and shows the difference in the file. The difference in archive indicates that there is a difference in one of its file, and the textual difference in that particular file can unveil the cause of unreproducibility. Thus, we reduce the search space to the most deeply nested unified_diff attribute. This also helps us to easily read the difference in the file.
We then write regular expressions to cluster the causes of unreproducibility based on file names and tool names reported in the source1 attribute and specific pattern in the unified_diff attribute. For example, if regex
@@
is matched, we infer that the reason of unreproducibility is due to timestamps in the file and then we manually inspect the difference in the file to validate the root cause of the unreproducibility. The manual inspection also helps us to identify more reasons for unreproducibility. For example, while looking through differences in timestamps, we also observe differences about version control system information. This iterative process through manual analysis continuously refines the taxonomy.
It is important to note that regular expressions are not directly mapped to the reasons we propose later, because this would have too many false positive. We rather rely on manual analysis to categorize differences. Thus, we do not count the causes of unreproducibility but build a comprehensive taxonomy about Java unreproducibility.
@@ 4.3. Results
Reason for Unreproducibility Root Cause of Unreproducibility Fine-grained Cause Novelty Main Mitigation
Build Manifest Environment Built-By Canonicalization by rebuilder
Signed JARs
Os-Version
Eclipse Properties
Rebuild Process Implementation-Build-Java-Vendor Fix rebuild process
Created-By
Originally-Created-By
SCM-Revision
Build-Jdk
Dynamic Properties SCM-Git-Branch Fix build process
Bnd-LastModified
Inconsistent Build Configuration Order of properties and their values Canonicalization by rebuilder
Software Bill of Materials Java Vendor Removal of hash algorithms Fix rebuild process
Inconsistent Build Configuration Addition of components Fix build process
Removal of components
Modification of components
Dynamic Properties Timestamp Canonicalization by rebuilder
SerialNumber
External Metadata License Canonicalization by rebuilder
Description of components
ExternalReferences
Filesystem Environment Permissions Canonicalization by rebuilder
Ownership
Size
Absolute Paths
Inconsistent Build Configuration Type Fix build process
Files removed or added
JVM Bytecode JDK Version Debug information Fix build process
Optimization and de-optimization
Refactoring
Embedded Metadata Absolute file paths, timestamps, Java/project version, Git properties, usernames Fix build process
Generated Code Java Compiler Canonicalization by rebuilder
Maven Plugins Fix rebuild process
Versioning Properties Source Repository State Number of commits Canonicalization by rebuilder
Number of Git tags
Branch name
Local branch name
Timezone of commit
Remote URL
Timestamps Build-time Variability Documentation Fix build process
Shell scripts
Executable binaries
JVM bytecode
Build manifest
File metadata
Table 2. The taxonomy of unreproducibility causes based on the analysis of 12,803 unreproducible artifacts. We are the first to study the most appropriate mitigation strategy.
In this section, we present a comprehensive taxonomy of unreproducibility in Java, based on our large-scale, systematic protocol presented in Figure 1. Table 2 summarizes the taxonomy. The first column lists the reason for unreproducibility. The second column lists the root cause that triggered the difference. The third column further breaks down the root cause into fine-grained causes. The fourth column indicates if the root cause is novel and has not been reported for Java artifacts in the literature before, including recent work by Xiong et al. (Xiong et al., 2022). Finally, the last column suggests the main strategy to fix the unreproducibility, one of the three presented in Subsection 2.3.
For example, consider the row where ‘JVM Bytecode’ is the reason and ‘Generated Code’ is the root cause. We break it down into two further causes - ‘Java Compiler’ and ‘Maven Plugins’. A build difference due to generated JVM bytecode by Maven plugins has not been reported in the literature before. However, a build difference due to Java compiler has been reported by Xiong et. al (Xiong et al., 2022).
For differences due to Java compiler, we can use jNorm which canonicalizes the differences. Differences due to Maven plugins can be mitigated by fixing the rebuild process.
@@ 4.3.1. Unreproducible Build Manifests
We consider two manifest files in our analysis — MANIFEST.MF and pom.properties. The MANIFEST.MF file is a mandatory metadata file in a JAR file. It potentially contains information on the entry point of the JAR file, signatures of the classes, and details of the version control system  [^13] https://docs.oracle.com/javase/tutorial/deployment/jar/manifestindex.html. pom.properties is an automatically generated file by Maven that contains the group ID, artifact ID, and version of the Maven module.
The first set of differences in manifests is due to a change in environment. For example, the Built-By attribute in the MANIFEST.MF file is sometimes updated to the username of the user who built the JAR file. Mitigation: This non-deterministic information should be stripped out from the JAR file by rebuilder before comparison. See section˜A.1, section˜A.4, section˜A.5, and section˜B.3 for more attributes that follow the same problem and solution.
The next set of differences in manifests is due to differences in the script that configures the rebuild. For example, the Build-Jdk attribute depends upon the version of JDK used to build the JAR file. Mitigation: All environment specifications must be exactly specified in the the build script by rebuilder. Otherwise, the difference is not only in the attribute but it impacts other components of the release too. For example, the Build-Jdk attribute in the MANIFEST.MF file is set to the version of JDK used to build the JAR file and the rebuilder should use the same version of JDK to rebuild the JAR file. As we will see later, a different JDK versions can lead to differences in the JVM bytecode and hence the artifact becomes unreproducible. See section˜A.2, section˜A.3, section˜A.6, section˜A.7, and section˜A.8 for concrete examples.
Finally, there are reasons of differences in the manifests MANIFEST.MF and pom.properties that are due to non-deterministic behavior in the build script of the builder. For example, build manifests pom.properties and MANIFEST.MF can differ in order of the attributes if they are not sorted. Mitigation: The builder should depend on plugins that generate sorted data  [^14] https://github.com/apache/maven-archiver/commit/763a940540eefad74f9ba73cb5eed288dc4e639d, should not embed non-deterministic data (e.g., branch names, timestamps) in the artifact, and stay up to date with the latest version of plugins. For example, only the latest version of Maven Archiver is able to maintain a deterministic file ordering. section˜A.9, section˜A.10, section˜A.11, section˜B.1, and section˜B.2 describe these reasons and solutions in detail.
@@ 4.3.2. Unreproducible Software Bill of Materials
The first set of changes are due to incorrect Java vendor used by the rebuilder. For example, old JDK versions do not have newer hash algorithms like SHA-3. However, vendors of some JDKs backport these algorithms and if builder uses a JDK with backported algorithms, the resulting SBOM will have hashes computed using more algorithms. If rebuilder does not use the same JDK, there will be differences in number of hash algorithms used. Mitigation: The rebuilder should use the same JDK vendor and version as the builder. This information may be present [^16] This is not a standard practice. These properties could either be embedded in MANIFEST.MF or module-info.java, or documented in the README of the source repository. Hence, this information may not always be available and its deduction can be adhoc. in the Build-Jdk and Implementation-Build-Java-Vendor attribute in the MANIFEST.MF file. See section˜C.1 for more details.
The next set of changes is due to a inconsistent build configuration by the builder. For example, the build process may use different ad-hoc scripts or rely on local caches to fetch components. Mitigation: The builder should use a conventional build process, such as only using standard Maven commands if only Maven projects are present in the project. Caches, like local Maven repository, should also be cleared because it can contain outdated components. This helps the rebuilder to reproduce the build process exactly without knowing all ad hoc scripting in the build. See section˜C.2 and section˜C.3 for concrete examples.
The third set of changes is due to volatile identifiers in the SBOM. CycloneDX attributes like timestamp and serialNumber are non-deterministic and hence they differ in each build. Mitigation: These attributes only act as an identifier and hence they can be canonicalized by the rebuilder before comparison. See section˜C.5 and section˜C.4 give concrete examples of releases falling into this case.
The last set of changes is due to external metadata in the SBOM. These attributes are fetched from external sources and hence if the external source updates metadata like license, description of component, or external references, the SBOM will be different for the subsequent builds. Mitigation: These three attributes can be canonicalized before comparison. In general, data from mutable external sources should be canonicalized. See section˜C.6, section˜C.7, and section˜C.8 which lists releases falling into this case.
We are the first ones to report that SBOM can be a reason of unreproducible builds.
@@ 4.3.3. Unreproducible Filesystem
We observe unreproducible Maven releases due to presence or absence of files, changes in the absolute file paths and their related metadata such as permissions, ownership, size and type. We classify them into two reasons below.
The first set of changes are due to an influence from the environment. For example, the permissions of the file can vary depending on the umask of the environment of rebuilder section˜D.2. The ownership and timestamp of the file can also vary depending who and when the artifact is built section˜D.3 section˜D.4. File sizes can vary due to newlines as operating systems handle it differently section˜D.6. The names of the files can vary if absolute paths are used to refer to the files in artifact section˜D.7. These names can be of generated files or they could be embedded in JVM bytecode and configuration files which makes the artifact unreproducible section˜D.8. Finally, the order of files in Jar file section˜G.5 can vary as the order depends upon the filesystem layout  [^17] https://reproducible-builds.org/docs/archives/. Mitigation: The rebuilder should either set a fixed value or strip this information from the metadata of the file. For example, the byte buffer of zip file can be edited to set the username and group of an entry in Jar file to an empty string.
Finally, there are some causes due to differences in the build process by the builder and rebuilder. For example, some binaries are not included in the rebuild version as they are not generated by the rebuilder build process. Mitigation: The rebuilder process should replicate the build process of the builder exactly. For the above example, the rebuilder should incorporate commands to generate the missing binaries. An automatic way to infer the correct process is to look for different Maven profiles and activate them during the rebuild process. See section˜D.1 and section˜D.5 for more concrete examples.
@@ 4.3.4. Unreproducible JVM bytecode
We observe 936 unreproducible Maven releases due to changes in the JVM bytecode when rebuilding. The reasons are as follows.
@@
The first set of changes are due to the use of different versions of the JDK or JDK flags used in the build and rebuild process. This differences are due to core compilation changes such as refactoring, ordering changes in the bytecode (specifically methods, fields, static initializers, entries in the constant pool, and array type values in annotation section˜G.1), optimization flags, debug flags. For example, java.nio.ByteBuffer.flip() returns a different type in JDK 8 and JDK 11. For io.dropwizard.metrics:metrics-collectd:4.1.20 rebuilt with JDK 11, all flip calls return java.nio.ByteBuffer which is different from the reference version built with JDK 8 where the return type is java.nio.Buffer.  [^18] This is reported in the bug tracking system of OpenJDK (Gafter, 2002) which shows that this change is introduced in JDK 9. See section˜E.1, section˜E.2, and section˜E.3 for more examples. Mitigation: The rebuilder should use the same JDK version and flags as the builder. We propose that the builder should always embed the precise vendor and version of JDK used to build the artifact in the MANIFEST file under the attribute Build-Jdk. This way, the rebuilder can use the same version and vendor of JDK to rebuild the artifact by reading the information instead of guessing it. If the builder is using a custom JDK as done in Google Guava (Moore, 2023), this should be made available to the rebuilder for verification.
@@
The next set of changes are due to embedding information related to environment. We observe that the bytecode has git branch names, user name of the builder, and absolute paths. For example, org.apache.hive:hive-exec:4.0.0-alpha-2 embeds git properties and timestamps in package-info.class file using a shell script that adds a Java annotation to package-info.java. The Java compiler also embeds the exact Java version used to compile the source code in the bytecode of module-info. However, this hinders reproducibility even if the patch version of the JDK is different. Mitigation: The builder should avoid embedding environment specific information in the bytecode. This information changes with each rebuilder and hence the bytecode will be different making the artifact unreproducible. See section˜E.4 for example of a release where it happens.
@@
Finally, the last set of changes are due to build time code generation or modification. The Java compiler produces synthetic code during build in order to handle access to type members (Rimenti, 2018). The compiler also names lambda functions (Evans, 2020) and anonymous classes (Maverik, 2016) which are not present in the source code. In addition to these capabilities, Maven provides plugins to generate or modify code during the build process. All of this generated code can be different across different builds and hence the build is unreproducible. For example, we observe a difference in names of lambda function in org.apache.helix:zookeeper-api:1.0.3. Another example is org.apache.hive:kafka-handler:4.0.0-alpha-2 which uses maven-shade-plugin to rename package names from kafkaesqueesque to kafkaesque but only when a Maven profile is activated. Mitigation: The names of the synthetic code can be canonicalized by the rebuilder before comparison. Bytecode canonicalization (Sharma et al., 2024) is a technique to strip out non-deterministic information from the bytecode. Maven can be configured to have multiple build profiles and this changes the artifacts produced. The rebuilder can infer the build profiles required to reproduce the build process by looking at the manifest file (pom.xml) of the artifact and choose the appropriate one which is designed for deployment. See section˜E.5 for more examples.